Find the position function so given the acceleration function and an initial value. 

a(t) = 12 - t, v(0) = 9, s(0) = 0


s(t) = _______ 
Answer #1

Solution given that all) = (2+; 10) =9, SCO) =0 u(t) = velocity - Salt)dt = 5 (12-1)dt V(t) = 12t- t2 tc → 0 when V()= > 16)